OpenStudy (anonymous):

will medal Marie mixes 16 liters of 78% acid solution with a 25% acid solution to make a 41% acid solution. How many liters of 25% solution did she use? 8 liters 12 liters 13.7 liters 37 liters

OpenStudy (fortytherapper):

Alright, so what I did was set up a proportion after calculating a % So first, would you know how to find out many liters were acid out of the 16? Basically, can you find what 78% of 16 is?

OpenStudy (fortytherapper):

To find out a percentage of a number, you take the number and multiply it by the decimal of the percentage over 1 So 78% would turn into .78, like you're just putting a dot in front of it. So you would do 16 times .78

OpenStudy (anonymous):

1248

OpenStudy (fortytherapper):

Yeah, but with the dot, it'll be 12.48 Now that we found this we can put it over 16.00, because that's the total amount of solution we have 12.48 ----- 16.00 Now they give you another total, which is 41% and they want to know how many liters it took to make the solution 25%. So first let's solve a proportion 12.48 25.00 ----- = ---- 16.00 x Would you know how to solve this? With cross multiplying and dividing?

OpenStudy (fortytherapper):

Sorry for the delay, I forgot how I solved it myself lol

OpenStudy (anonymous):

yea no I don't know how to solve that

OpenStudy (anonymous):

is the second part like the first part where I have to times it?

OpenStudy (fortytherapper):

Yes, but you would do it to the 25 and 16 since they are "crossing" each other. After you do that, you would divide that by 12.48, giving you X

OpenStudy (anonymous):

? I got 32.05 16 x 25 = 400 (even when it is 16.00 and 25.00) and then divide it by 12.48 and I got 32.05?

OpenStudy (welshfella):

another approach would be to build an equation:- work in wieghts of pure acid 16*0.78 + 0.25x = 0.41( 16 + x) where x is the volume of the 25% acid

OpenStudy (fortytherapper):

Yep. The 32.05 is 100% solution, but they asked for 25% solution. So what would you do? Omg I like welsh's approach. I really should learn now to over complicate things >__>

OpenStudy (welshfella):

Yea - try to make it as simple as possible.

OpenStudy (welshfella):

16*0.78 + 0.25x = 0.41( 16 + x) 12.48 + 0.25x = 0.41x + 6.56 5.92 = 0.16x x = 5.92/0.16

OpenStudy (welshfella):

37 liters